### v1

# Quantitative Measures for Passive Sonar Texture Analysis

Abstract
Passive sonar signals contain complex characteristics often arising from environmental noise, vessel machinery, and propagation effects. While convolutional neural networks (CNNs) perform well on passive sonar classification tasks, they can struggle with statistical variations that occur in the data. To investigate this limitation, synthetic underwater acoustic datasets are generated that centered on amplitude and period variations. Two metrics are proposed to quantify and validate these characteristics in the context of statistical and structural texture for passive sonar. These measures are applied to real-world passive sonar datasets to assess texture information in the signals and correlate the performances of the models. Results show that CNNs underperform on statistically textured signals, but incorporating explicit statistical texture modeling yields consistent improvements. These findings highlight the importance of quantifying texture information for passive sonar classification. Code is publicly available: https://github.com/Advanced-Vision-and-Learning-Lab/Passive_Sonar_Texture_Analysis.
1 Introduction
Texture is a fundamental property in both image and audio domains, capturing the underlying structural and statistical patterns in signals [1]. In images, texture refers to spatial elements such as edges, repetitive structures, and intensity variations. In the audio domain, structural texture corresponds to periodic features like harmonics and modulations [2], while statistical texture reflects stochastic variations in amplitude and frequency [3]. These texture characteristics influence how models interpret signals, especially for complex tasks like passive sonar classification. In underwater sensing, variations in passive sonar may arise due to environmental noise, turbulence, and reflections from dynamic sources [4]. Conventional features such as Mel-frequency cepstral coefficients (MFCCs), Gammatone filters, and wavelet transforms are widely used. However, they often fail to capture how statistical randomness interacts with structured periodicity [5, 6].
To represent texture more effectively, prior work has explored statistical descriptors including mean, variance, skewness, entropy, and other higher-order distributional measures [7, 8]. These help distinguish between noise-like patterns and harmonically rich signals [9]. Structural texture has been examined using methods such as autocorrelation and harmonic envelope modeling, which capture temporal repetition and spectral regularity [10]. Despite the importance of both texture types, most deep learning models, especially convolutional neural networks, are optimized for capturing local structural features [8]. As a result, they may struggle with capturing statistical texture features [11, 12].
The Histogram Layer Time-Delay Neural Network (HLTDNN) [13], a particular type of CNN demonstrated that incorporating statistical representations improves passive sonar classification. Building on this work, this manuscript explores whether explicitly quantifying audio texture can provide more guidance on model selection for passive sonar classification. Synthetic datasets are developed to investigate new metrics. Specifically, texture characteristics are validated using two quantification metrics: one entropy-based for statistical texture and another autocorrelation-based for structural texture. These metrics are subsequently applied to real-world datasets to examine how texture composition relates to model performance.
The main contributions of this work are as follows:
  • Synthetic underwater acoustic dataset with controlled statistical and structural textures variations.
  • Formal definitions to quantify passive sonar texture.
  • Empirical analysis demonstrating the correlation between texture characteristics of the data and model performance.
2 Methodology
Three synthetic datasets are designed to simulate sonar-like recordings with controlled texture characteristics that focus on variations of (1) amplitude (statistical texture), (2) period (structural texture), and (3) both amplitude and period (mixed statistical and structural textures). Each dataset includes four vessel classes with 10,000 samples per class, resulting in 40,000 samples per dataset. The signal design is guided by observations from the DeepShip dataset [14], where cargo and tanker vessels tend to exhibit lower-frequency signatures, while passengership and tug vessels show greater activity in higher-frequency bands. Reflecting this, two vessel classes in each subset are assigned lower-frequency bands and the other two higher-frequency bands. Figure 1 compares spectrograms across the three synthetic datasets. The generation of each subset is detailed in Sections 2.1 through 2.3.
Figure 1: Example spectrograms showing one sample per class (Vessel 1, Vessel 2, Vessel 3, Vessel 4) for each synthetic dataset type: Structural Texture, Statistical Texture, and Mixed Statistical and Structural Textures.
2.1 Statistical Texture Dataset
This dataset emphasizes stochastic variation in amplitude and background noise while introducing partial spectral overlap to simulate subtle structural cues. Each ship class is assigned three frequency components, with one shared across adjacent classes to mimic real-world spectral overlap. For instance, Vessel 1 spans 3000–5500 Hz while Vessel 2 spans 4000–5500 Hz; similar overlap continues across all four classes. Each signal is segmented into four equal-duration regions. Within each segment, the amplitude envelope $A^{f}(t)$, defined for each frequency component $f$ over time $t\in[0,T]$, is randomly sampled from either a Rayleigh or K-distribution [3] as shown in 1:
$A^{f}(t)\sim\begin{cases}\text{Rayleigh}(\sigma),&\text{with probability }p,\\ \text{K}(k,\theta),&\text{with probability }1-p,\end{cases}$ (1)
where $\sigma=0.5$ controls the spread of the Rayleigh distribution, while $k=14$ and $\theta=1.0$ define the shape and scale of the K-distribution. These segment-wise transitions are smoothed using sinusoidal amplitude modulation to avoid abrupt energy shifts. Each class uses a unique combination of modulation depth and rate to vary temporal dynamics.
In 2, background noise $\eta(t)$ is added to each signal that consists of class-specific colored noise $\nu(t)$ and time impulses:
$\eta(t)=\nu(t)+\sum_{k\in\mathcal{I}}a_{k}\cdot\delta(t-t_{k}),$ (2)
where $\nu(t)$ represents white or pink noise depending on the class, $\mathcal{I}$ is a random set of impulse time indices, and $a_{k}\sim\mathcal{U}[-1,1]$ determines the impulse amplitude. For example, Vessel 1 combines white and pink noise, Vessel 2 uses pink noise only, Vessel 3 adds short impulses over white noise, and Vessel 4 incorporates impulsive pink noise. This configuration enriches temporal and spectral variability while preserving distinguishable class characteristics.
2.2 Structural Texture Dataset
This dataset emphasizes periodic and deterministic temporal structure with minimal randomness. Each signal consists of three harmonics derived from a class-specific base frequency $f_{c}$, with slight jitter applied to introduce spectral variation as defined in 3:
$f_{k}=k\cdot f_{c}+\delta_{k},\quad\delta_{k}\sim\mathcal{U}[-50,50],$ (3)
where $f_{k}$ is the frequency of the $k$-th harmonic, and $\delta_{k}$ introduces small deviations (sampled from a uniform distribution) to mimic natural variability. Each vessel class is assigned a unique base frequency and amplitude envelope shape to produce distinct structural patterns. Vessels 1 through 4 use base frequencies of 2800 Hz, 3600 Hz, 5400 Hz, and 6200 Hz, respectively, paired with triangular, exponential decay, ramp-up, and ramp-down envelopes. These deterministic envelopes, defined over normalized time $t\in[0,1]$, promote strong periodicity and reduce intra-class randomness. Low-amplitude zero-mean Gaussian noise is added to each signal to not mask the harmonic structure. The standard deviation for the noise was randomly sampled from the range $[0.006,0.01]$. The overall design ensures that each class exhibits strong structural regularity while preserving slight temporal and spectral variability, enabling the evaluation of model sensitivity to well-defined periodic features.
2.3 Mixed Statistical and Structural Textures Dataset
This dataset is designed to simulate acoustic signals that simultaneously exhibit variations in amplitude and period. This combination reflects more realistic underwater sounds where both deterministic patterns and random fluctuations coexist. Each amplitude envelope $\hat{A}^{f}(t)$ is generated by modulating a probabilistic envelope with a sinusoidal component as shown in 4:
$\hat{A}^{f}(t)=\alpha(t)\cdot\left(1+d\cdot\sin(2\pi rt)\right),$ (4)
where $\alpha(t)\sim\text{Rayleigh}(\sigma)$ or $\text{K}(k,\theta)$ represents the statistical envelope shape, and the sinusoidal term introduces periodic structure. The modulation depth $d$ and rate $r$ are sampled from class-specific uniform ranges to emulate realistic fluctuations and vessel-dependent vibration patterns.
To ensure smooth temporal evolution between statistical textures, a Gaussian-based blending function is applied to transition between Rayleigh and K-distributed envelopes over the full signal duration. Specifically, for each time point $t$, a weight function $w(t)$ based on a normalized Gaussian kernel is used to compute a blended amplitude envelope using 5:
$\alpha(t)=w(t)\cdot\alpha_{\text{Rayleigh}}(t)+\left(1-w(t)\right)\cdot\alpha_{\text{K}}(t),$ (5)
where $w(t)=\exp\left(-\frac{(t-\mu)^{2}}{2\sigma^{2}}\right)$ controls the influence of each distribution at time $t$, and $\mu,\sigma$ are class-specific parameters determining the center and width of the texture transition.
Each class uses a unique combination of frequency bands, amplitude distribution sequences, and modulation properties. Vessel classes with lower-frequency profiles (Vessel 1 and Vessel 2) are assigned slower amplitude modulations, reflecting the slower mechanical oscillations of large ships such as cargo or tankers. In contrast, higher-frequency classes (Vessel 3 and Vessel 4) simulate smaller, faster vessels that typically exhibit quicker modulation patterns due to faster propeller activity and particular structural dynamics [15]. Table 1 summarizes the design choices for the four vessel classes.
Table 1: Class-wise configuration for the mixed statistical and structural textures dataset.
Vessel Class Frequencies (Hz) Amplitude Distributions Modulation Rate
[…]
2.4 Statistical Texture Quantification via Bidirectional Temporal Entropy
To quantify statistical texture in passive sonar signals, the bidirectional entropy-based method inspired by the temporal texturedness indicator proposed in [16] is adopted for this work. This approach evaluates how Shannon entropy accumulates over time in both forward and reverse directions of a signal. The underlying premise is that statistically textured signals exhibit consistent randomness over time, leading to minimal asymmetry between the two entropy trajectories. Let $x(t)$ be the input signal defined over time $t\in[0,T_{\text{obs}}]$, where $T_{\text{obs}}$ is the total observation duration. The signal is sampled at a frequency $f_{s}$ and divided into overlapping frames of length $L$ with hop size $H$. For each frame index $i$, the signal’s amplitude values are binned into a histogram with $M$ total bins. The resulting histogram is then normalized to obtain the probability distribution $\{p_{i,k}\}$, where $p_{i,k}$ denotes the probability corresponding to the $k$-th bin. The entropy at each step is computed using 6:
$H_{i}=-\sum_{k=1}^{M}p_{i,k}\log_{2}p_{i,k}.$ (6)
Entropy is accumulated in two directions: forward using the original signal $x(t)$, and reverse using the time-inverted signal $x_{r}(t)=x(T_{\text{obs}}-t)$. This results in two cumulative entropy curves for the forward and reverse directions, denoted $H_{d}(i)$ and $H_{r}(i)$, respectively.
The surface area $S$ between these curves captures directional entropy asymmetry and is normalized by the theoretical maximum $S_{\text{max}}=T_{\text{obs}}\cdot\log_{2}M$, which corresponds to the entropy of a white noise signal. The relative entropy is defined as $P_{\text{rel}}=\frac{H_{\text{final}}}{\log_{2}M}$, where $H_{\text{final}}$ is the final cumulative entropy in the forward direction. The variable $P_{T}$ represents the proportion of the signal duration required for the forward and reverse entropy curves to converge. Using these components, the final Statistical Texture Score (StaTS) is defined in 7:
$\text{StaTS}_{\text{stat}}=5\cdot P_{\text{rel}}\cdot\left(1-P_{T}\cdot\frac{S}{S_{\text{max}}}\right),$ (7)
Higher scores reflect signals with strong statistical texture, characterized by high entropy, rapid forward–reverse convergence, and minimal directional asymmetry.
2.5 Structural Texture Quantification via Autocorrelation
The Structural Texture Score (StrTS) is introduced to quantify periodicity and structural regularity in passive sonar signals using the normalized autocorrelation function. Autocorrelation measures the similarity between a signal and its time-shifted versions, making it effective for identifying repeating harmonic or modulated patterns [17]. The signal is first transformed to be zero-mean before computing the autocorrelation. To ensure robustness across signals of different durations, the maximum lag is set adaptively as 25% of the total signal duration, with safeguards to prevent it from being too short or excessively long. This approach ensures that the method captures relevant structural patterns without being affected by noise or irrelevant long-term dependencies. Only positive lags beyond zero are considered, and the first significant positive peak is selected as the dominant periodic component. Negative peaks, indicating anti-phase alignment, are excluded.
Let $R(\tau)$ denote the normalized autocorrelation value at lag $\tau$, where $\tau\in[1,\tau_{\text{max}}]$, and $\tau_{\text{max}}$ is the maximum lag considered. The lag corresponding to the first significant positive peak is identified as the dominant periodic component as shown in 8:
$\tau^{*}=\arg\max_{\tau\in[1,\tau_{\text{max}}]}R(\tau),$ (8)
The corresponding autocorrelation value at this lag, $R(\tau^{*})$, reflects the strength of the periodic structure and is scaled to produce the StrT in 9:
$\text{StrTS}=5\cdot R(\tau^{*}),$ (9)
The score lies within the range $[0,5]$, where higher values indicate stronger and more regular periodicity. For consistency, both StrTS and StaTS share this 0–5 scale, where lower scores indicate weak periodicity or less texture and higher scores reflect stronger structural or statistical patterns.
In addition to the scalar score, self-similarity matrices offer a complementary visual representation of structural repetition. These matrices capture pairwise similarity between overlapping frames of the signal, highlighting the repeated temporal pattern. Strong diagonal structures in the matrix correspond to high self-similarity, while noisy or unstructured matrices indicate a weak or lack of repetition.
Figure 2: Self-similarity matrices for three representative signals. (a) White noise (StrTS = 0.05): lacks clear structure. (b) Cargo ship signal (StrTS = 2.73): shows partial diagonal patterns. (c) Pure sine wave (StrTS = 5.00): exhibits strong and evenly spaced diagonals.
As shown in Fig. 2, these visualizations qualitatively reflect periodic structure and align well with the StrTS scores. Higher StrTS scores correspond to clearer, more structured diagonal patterns, reinforcing the metric’s interpretability and relevance as a measure of structural texture.
3 Experimental Procedure
3.1 Dataset Description
Table 2: Dataset summary with number of classes and sample counts.
Dataset Type Classes Total Samples
Synthetic-Structural Synthetic 4 40,000
Synthetic-Statistical Synthetic 4 40,000
Synthetic-Balanced Synthetic 4 40,000
DeepShip [14] Real-world 4 33,770
VTUAD [18] Real-world 5 175,965
Experiments are conducted on three synthetic and two real-world datasets to evaluate model performance under varying audio texture conditions. The real-world datasets DeepShip [14] and Vessel Type Underwater Acoustic Data (VTUAD) [18] offer diverse passive sonar recordings. VTUAD is the largest with 175,965 samples across five classes. DeepShip provides 33,770 labeled samples across four classes. Table 2 summarizes the datasets by type, number of classes, and sample count.
3.2 Experimental Design
Table 3: Comparison of structural (StrTS) and statistical (StaTS) texture scores along with classification accuracies of TDNN[19] and HLTDNN [13] across five datasets. The last column shows the absolute accuracy improvement ($\Delta$ Acc) of HLTDNN over TDNN. The best average accuracy is bolded.
Dataset StrTS StaTS TDNN Acc. (%) HLTDNN Acc. (%) $\Delta$ Acc. (%)
Synthetic-Structural $4.96\pm 0.03$ $3.52\pm 0.09$ $72.50\pm 3.38$ $99.67\pm 0.15$ +27.17
Synthetic-Statistical $3.76\pm 0.35$ $4.93\pm 0.04$ $34.85\pm 1.04$ $55.05\pm 1.13$ +20.20
Synthetic-Mixed $4.39\pm 0.21$ $4.33\pm 0.56$ $72.18\pm 3.21$ $99.08\pm 0.88$ +26.90
DeepShip $3.16\pm 0.64$ $4.43\pm 0.55$ $50.49\pm 0.74$ $59.62\pm 1.69$ +8.06
VTUAD $4.45\pm 0.16$ $4.76\pm 0.32$ $71.92\pm 2.86$ $83.45\pm 0.57$ +10.53
All audio signals are resampled to 32 kHz for consistency. Each resampled signal is segmented into five-second intervals, except for the VTUAD dataset, which provides pre-segmented one-second clips with predefined dataset partitions. Segments are converted into logarithmic mel-frequency spectrograms using a Hanning window (window size 1024, hop 320) with 1024 mel-frequency bins. No data augmentation is applied to preserve the native texture characteristics. For all datasets (except VTUAD), samples are split into 70% training, 10% validation, and 20% test sets. Classification experiments use two models: TDNN as the baseline, and HLTDNN, which incorporates a histogram-based feature layer to enhance statisical texture representation. Both models are trained for 100 epochs using the Adagrad optimizer (learning rate $1\times 10^{-3}$, batch size 128). Test set accuracy is used for model evaluation. In addition to model performance, each dataset is analyzed using the two texture metrics (StrTS and StaTS) to provide insight into the texture properties of the signals.
4 Results and Discussion
4.1 Synthetic and Real-World Datasets Evaluation
Table 3 summarizes the structural (StrTS) and statistical (StaTS) texture scores, along with classification accuracies for TDNN and HLTDNN across five datasets. The last column reports the accuracy improvement ($\Delta$ Accuracy) of HLTDNN over TDNN. The results support the hypothesis that HLTDNN, with its histogram-based representation, is better suited for statistically rich signals. In the statistical texture dataset, where StaTS is highest ($4.93\pm 0.04$), HLTDNN achieves a substantial +20.20% improvement over TDNN. This validates HLTDNN’s capacity to capture statistical variations more effectively than TDNN. The structural texture dataset exhibits the highest StrTS ($4.96\pm 0.03$) and a lower StaTS. While TDNN performs reasonably well ($72.50\pm 3.38$), HLTDNN still outperforms it with a +27.17% gain, demonstrating that its statistical modeling does not hinder its ability to capture structural patterns.
In the mixed statistical and structural textures dataset, where both texture types coexist, HLTDNN achieves near-perfect accuracy ($99.08\pm 0.88$) and shows a +26.90% gain over TDNN. This highlights HLTDNN’s adaptability in handling hybrid texture profiles. On the DeepShip dataset, which exhibits moderate statistical and lower structural texture complexity, HLTDNN outperforms TDNN by +8.06%. Despite being a real-world dataset with inherent variability, the gain shows HLTDNN’s effectiveness in generalizing to practical scenarios. VTUAD, the largest dataset, contains high values in both StrTS and StatS, and HLTDNN again shows a strong improvement of +10.53%. This result confirms the model’s scalability and ability to learn from complex, large-scale real-world data.
4.2 Pretraining Models Using Synthetic Datasets
Table 4: Model performance (%) on DeepShip with and without synthetic mixed texture pretraining. The best average accuracy is bolded.
Model Without Pretraining With Pretraining
TDNN $50.49\pm 0.74$ $52.5\pm 1.29$
HLTDNN $59.62\pm 1.69$ $62.11\pm 1.36$
To further investigate texture-aware learning, both models were pretrained on the mixed statistical and structural textures dataset and fine-tuned on DeepShip since the synthetic dataset was modeled after DeepShip’s characteristics. The performance comparison is shown in Table 4. Pretraining led to a consistent gain for both models, with HLTDNN improving by +2.49%. These results suggest that synthetic pretraining is particularly effective when the texture characteristics of the synthetic and real-world domains are well aligned. In summary, HLTDNN consistently outperforms TDNN across all datasets. The gains are especially notable in statistically textured conditions and further enhanced by texture-aligned pretraining. These findings highlight the benefit of explicitly modeling statistical texture and tailoring synthetic data to match real-world texture profiles.
5 Conclusion
This work investigated the role of structural and statistical texture in CNN-based passive sonar classification. Statistically rich signals frequently occur in passive sonar environments due to stochastic propagation, scattering, and background interference — phenomena that standard CNNs may fail to capture effectively. By constructing synthetic underwater datasets with controlled texture attributes, it was possible to isolate and analyze model behavior across varying degrees of signal complexity. Experimental findings show that traditional CNN architectures such as TDNN are more sensitive to structural regularities, whereas the texture-aware HLTDNN demonstrates improved performance on signals characterized by strong statistical texture.
These results support the importance of explicitly modeling statistical texture in passive acoustic target classification. The proposed statistical and structural texture quantification metrics further contribute to model interpretability by linking signal properties to performance characteristics. Future directions include generating synthetic data samples to adapt audio foundation models such as AudioPaLM [20] and AudioGPT [21] for passive sonar, analyzing texture sensitivity in transformer-based architectures such as the Audio Spectrogram Transformer (AST) [22], embedding texture-aware regularization into training objectives, leveraging more advanced approaches for data generation such as diffusion models [23] and applying the texture metrics to other audio domains.
6 Acknowledgment
Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.

### v2

@@
-Quantitative Measures for Passive Sonar Texture Analysis
+Further author information: (Send correspondence to J.P.)
 Abstract
@@
 These findings highlight the importance of quantifying texture information for passive sonar classification.
-Code is publicly available: https://github.com/Advanced-Vision-and-Learning-Lab/Passive_Sonar_Texture_Analysis.
+The code is publicly available at: https://github.com/Advanced-Vision-and-Learning-Lab/Passive_Sonar_Texture_Analysis.
+keywords:
+Passive sonar, statistical texture, structural texture, synthetic dataset, texture analysis, texture quantification.
+00footnotetext: DISTRIBUTION STATEMENT A.
+Approved for public release.
+Distribution is unlimited.
+This material is based upon work supported by the Under Secretary of War for Research and Engineering under Air Force Contract No.
+FA8702-15-D-0001 or FA8702-25-D-B002.
+Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Under Secretary of Defense for Research and Engineering.
+© 2025 Massachusetts Institute of Technology.
+Delivered to the U.S.
+Government with Unlimited Rights, as defined in DFARS Part 252.227-7013 or 7014 (Feb 2014).
+Notwithstanding any copyright notice, U.S.
+Government rights in this work are defined by DFARS 252.227-7013 or DFARS 252.227-7014 as detailed above.
+Use of this work other than as specifically authorized by the U.S.
+Government may violate any copyrights that exist in this work.
 1 Introduction
@@
 In the audio domain, structural texture corresponds to periodic features like harmonics and modulations [21], while statistical texture reflects stochastic variations in amplitude and frequency [23].
-These texture characteristics influence how models interpret signals, especially for complex tasks like passive sonar classification.
-In underwater sensing, variations in passive sonar may arise due to environmental noise, turbulence, and reflections from dynamic sources [4].
-Conventional features such as Mel-frequency cepstral coefficients (MFCCs), Gammatone filters, and wavelet transforms are widely used.
-However, they often fail to capture how statistical randomness interacts with structured periodicity [5, 6].
-To represent texture more effectively, prior work has explored statistical descriptors including mean, variance, skewness, entropy, and other higher-order distributional measures [7, 8].
-These help distinguish between noise-like patterns and harmonically rich signals [9].
+These characteristics influence how models interpret signals, especially for complex tasks like passive sonar classification.
+Variations in passive sonar arise from environmental noise, turbulence, and reflections from dynamic sources [15].
+While conventional features such as MFCCs, Gammatone filters, and wavelet transforms are widely used, they often fail to capture how statistical randomness interacts with structured periodicity [1, 13].
+Prior work has explored statistical descriptors including mean, variance, skewness, entropy, and other higher-order measures [7, 17] to distinguish between noise-like and harmonically rich signals [3].
 Structural texture has been examined using methods such as autocorrelation and harmonic envelope modeling, which capture temporal repetition and spectral regularity [22].
-Despite the importance of both texture types, most deep learning models, especially convolutional neural networks, are optimized for capturing local structural features [8].
-As a result, they may struggle with capturing statistical texture features [11, 12].
-The Histogram Layer Time-Delay Neural Network (HLTDNN) [13], a particular type of CNN demonstrated that incorporating statistical representations improves passive sonar classification.
-Building on this work, this manuscript explores whether explicitly quantifying audio texture can provide more guidance on model selection for passive sonar classification.
-Synthetic datasets are developed to investigate new metrics.
-Specifically, texture characteristics are validated using two quantification metrics: one entropy-based for statistical texture and another autocorrelation-based for structural texture.
-These metrics are subsequently applied to real-world datasets to examine how texture composition relates to model performance.
+Despite their importance, most deep models like CNNs are primarily optimized for local structural patterns [17], and often underperform in capturing statistical texture [18, 10].
+The Histogram Layer Time-Delay Neural Network (HLTDNN) [19] demonstrated that incorporating statistical representations improves passive sonar classification.
+Building on this, this work investigates whether explicitly quantifying audio texture can better guide model selection.
+Synthetic datasets are developed to validate two texture quantification metrics: one entropy-based for statistical texture and another autocorrelation-based for structural texture.
+These are also applied to real datasets to analyze how texture composition relates to performance.
 The main contributions of this work are as follows:
@@
 2 Methodology
-Three synthetic datasets are designed to simulate sonar-like recordings with controlled texture characteristics that focus on variations of (1) amplitude (statistical texture), (2) period (structural texture), and (3) both amplitude and period (mixed statistical and structural textures).
-Each dataset includes four vessel classes with 10,000 samples per class, resulting in 40,000 samples per dataset.
-The signal design is guided by observations from the DeepShip dataset [14], where cargo and tanker vessels tend to exhibit lower-frequency signatures, while passengership and tug vessels show greater activity in higher-frequency bands.
-Reflecting this, two vessel classes in each subset are assigned lower-frequency bands and the other two higher-frequency bands.
-Figure 1 compares spectrograms across the three synthetic datasets.
-The generation of each subset is detailed in Sections 2.1 through 2.3.
+Three synthetic datasets were created to simulate sonar-like recordings with controlled texture characteristics: (1) statistical texture based on amplitude variation, (2) structural texture based on periodic patterns, and (3) mixed statistical and structural textures.
+Each dataset contains four vessel classes with 10,000 samples per class, totaling 40,000 samples.
+The design was informed by patterns observed in the DeepShip dataset [9], where cargo and tanker vessels show stronger low-frequency signatures, and passengership and tug vessels show more high-frequency activity.
+To reflect this, the synthetic datasets assign lower frequencies to Classes 1 and 2 and higher frequencies to Classes 3 and 4.
+Spectral peaks were visually inspected using STFT plots of DeepShip recordings, where bright, stable regions in the frequency axis were noted across multiple samples per class.
+Larger vessels consistently exhibited high-energy content between 2.5–4.5 kHz, while smaller vessels showed dominant peaks between 5.0–6.5 kHz, guiding the selection of class-wise base frequencies.
+Amplitude distributions were modeled using Rayleigh and K-distributions, common in sonar signal modeling [23], with parameters empirically chosen to provide diverse yet realistic signal textures.
+Amplitude modulation depth and rate were sampled from ranges of [0.1, 0.3] and [1, 3] Hz, and noise levels from $[0.006,0.01]$, ensuring that background effects do not overwhelm key signal characteristics.
+As summarized in Table 1, each vessel class is configured with distinct frequency, amplitude, and modulation properties.
+Figure 1 shows representative spectrograms from each dataset.
+The full dataset generation procedures are described in Sections 2.1 through 2.3.
 Figure 1: Example spectrograms showing one sample per class (Vessel 1, Vessel 2, Vessel 3, Vessel 4) for each synthetic dataset type: Structural Texture, Statistical Texture, and Mixed Statistical and Structural Textures.
+Table 1: Class-wise configuration for the mixed statistical and structural textures dataset. Frequencies are expressed in kilohertz (kHz).
+Vessel Class Frequencies (kHz) Amplitude Distributions Modulation Rate
+Vessel 1 {2.2, 3.5, 5.0, 6.5} K $\rightarrow$ Rayleigh Slow
+Vessel 2 {3.0, 4.5, 6.0, 7.0} Rayleigh $\rightarrow$ K Slow
+Vessel 3 {9.0, 10.5, 12.0, 13.0} Rayleigh $\rightarrow$ K Fast
+Vessel 4 {10.0, 11.5, 13.0, 14.0} K $\rightarrow$ Rayleigh Fast
 2.1 Statistical Texture Dataset
-This dataset emphasizes stochastic variation in amplitude and background noise while introducing partial spectral overlap to simulate subtle structural cues.
-Each ship class is assigned three frequency components, with one shared across adjacent classes to mimic real-world spectral overlap.
-For instance, Vessel 1 spans 3000–5500 Hz while Vessel 2 spans 4000–5500 Hz; similar overlap continues across all four classes.
-Each signal is segmented into four equal-duration regions.
-Within each segment, the amplitude envelope $A^{f}(t)$, defined for each frequency component $f$ over time $t\in[0,T]$, is randomly sampled from either a Rayleigh or K-distribution [3] as shown in 1:
-$A^{f}(t)\sim\begin{cases}\text{Rayleigh}(\sigma),&\text{with probability }p,\\ \text{K}(k,\theta),&\text{with probability }1-p,\end{cases}$ (1)
-where $\sigma=0.5$ controls the spread of the Rayleigh distribution, while $k=14$ and $\theta=1.0$ define the shape and scale of the K-distribution.
-These segment-wise transitions are smoothed using sinusoidal amplitude modulation to avoid abrupt energy shifts.
-Each class uses a unique combination of modulation depth and rate to vary temporal dynamics.
-In 2, background noise $\eta(t)$ is added to each signal that consists of class-specific colored noise $\nu(t)$ and time impulses:
+This dataset emphasizes stochastic variation in amplitude and background noise, with partial spectral overlap between adjacent classes.
+Each signal contains three frequency components per class.
+The amplitude envelope $A^{f}(t)$, defined for each frequency $f$, is sampled as:
+$A^{f}(t)\sim\begin{cases}\text{Rayleigh}(\sigma),&\text{with probability }p,\\ \text{K}(k,\theta),&\text{with probability }1-p\end{cases}$ (1)
+To ensure smooth transitions, sinusoidal amplitude modulation is applied segment-wise. Background noise $\eta(t)$ is added as:
 $\eta(t)=\nu(t)+\sum_{k\in\mathcal{I}}a_{k}\cdot\delta(t-t_{k}),$ (2)
-where $\nu(t)$ represents white or pink noise depending on the class, $\mathcal{I}$ is a random set of impulse time indices, and $a_{k}\sim\mathcal{U}[-1,1]$ determines the impulse amplitude.
-For example, Vessel 1 combines white and pink noise, Vessel 2 uses pink noise only, Vessel 3 adds short impulses over white noise, and Vessel 4 incorporates impulsive pink noise.
-This configuration enriches temporal and spectral variability while preserving distinguishable class characteristics.
+where $\nu(t)$ is colored noise and impulses $a_{k}$ simulate short acoustic bursts. The design maintains amplitude-driven variation while limiting periodic structure.
 2.2 Structural Texture Dataset
-This dataset emphasizes periodic and deterministic temporal structure with minimal randomness.
-Each signal consists of three harmonics derived from a class-specific base frequency $f_{c}$, with slight jitter applied to introduce spectral variation as defined in 3:
+This dataset emphasizes deterministic temporal structure with minimal randomness.
+Each signal consists of three harmonics derived from a class-specific base frequency $f_{c}$, perturbed by uniformly sampled jitter:
 $f_{k}=k\cdot f_{c}+\delta_{k},\quad\delta_{k}\sim\mathcal{U}[-50,50],$ (3)
-where $f_{k}$ is the frequency of the $k$-th harmonic, and $\delta_{k}$ introduces small deviations (sampled from a uniform distribution) to mimic natural variability.
-Each vessel class is assigned a unique base frequency and amplitude envelope shape to produce distinct structural patterns.
-Vessels 1 through 4 use base frequencies of 2800 Hz, 3600 Hz, 5400 Hz, and 6200 Hz, respectively, paired with triangular, exponential decay, ramp-up, and ramp-down envelopes.
-These deterministic envelopes, defined over normalized time $t\in[0,1]$, promote strong periodicity and reduce intra-class randomness.
-Low-amplitude zero-mean Gaussian noise is added to each signal to not mask the harmonic structure.
-The standard deviation for the noise was randomly sampled from the range $[0.006,0.01]$.
-The overall design ensures that each class exhibits strong structural regularity while preserving slight temporal and spectral variability, enabling the evaluation of model sensitivity to well-defined periodic features.
+Each class uses a distinct amplitude envelope (e.g., triangular, exponential decay) defined over normalized time.
+Low-amplitude Gaussian noise is added to preserve harmonic clarity while introducing mild background variation.
 2.3 Mixed Statistical and Structural Textures Dataset
-This dataset is designed to simulate acoustic signals that simultaneously exhibit variations in amplitude and period.
-This combination reflects more realistic underwater sounds where both deterministic patterns and random fluctuations coexist.
-Each amplitude envelope $\hat{A}^{f}(t)$ is generated by modulating a probabilistic envelope with a sinusoidal component as shown in 4:
+This dataset simulates acoustic signals exhibiting both amplitude variation and periodic structure, capturing realistic underwater texture characteristics where random fluctuations and deterministic patterns coexist.
+Each amplitude envelope $\hat{A}^{f}(t)$ is generated by modulating a probabilistic envelope with a sinusoidal component:
 $\hat{A}^{f}(t)=\alpha(t)\cdot\left(1+d\cdot\sin(2\pi rt)\right),$ (4)
-where $\alpha(t)\sim\text{Rayleigh}(\sigma)$ or $\text{K}(k,\theta)$ represents the statistical envelope shape, and the sinusoidal term introduces periodic structure.
-The modulation depth $d$ and rate $r$ are sampled from class-specific uniform ranges to emulate realistic fluctuations and vessel-dependent vibration patterns.
-To ensure smooth temporal evolution between statistical textures, a Gaussian-based blending function is applied to transition between Rayleigh and K-distributed envelopes over the full signal duration.
-Specifically, for each time point $t$, a weight function $w(t)$ based on a normalized Gaussian kernel is used to compute a blended amplitude envelope using 5:
+where $\alpha(t)\sim\text{Rayleigh}(\sigma)$ or $\text{K}(k,\theta)$, and $d$, $r$ are class-specific modulation depth and rate parameters.
+To enable smooth transitions between amplitude distributions, a Gaussian-based blending function interpolates between Rayleigh and K-distributed envelopes:
 $\alpha(t)=w(t)\cdot\alpha_{\text{Rayleigh}}(t)+\left(1-w(t)\right)\cdot\alpha_{\text{K}}(t),$ (5)
-where $w(t)=\exp\left(-\frac{(t-\mu)^{2}}{2\sigma^{2}}\right)$ controls the influence of each distribution at time $t$, and $\mu,\sigma$ are class-specific parameters determining the center and width of the texture transition.
-Each class uses a unique combination of frequency bands, amplitude distribution sequences, and modulation properties.
-Vessel classes with lower-frequency profiles (Vessel 1 and Vessel 2) are assigned slower amplitude modulations, reflecting the slower mechanical oscillations of large ships such as cargo or tankers.
-In contrast, higher-frequency classes (Vessel 3 and Vessel 4) simulate smaller, faster vessels that typically exhibit quicker modulation patterns due to faster propeller activity and particular structural dynamics [15].
-Table 1 summarizes the design choices for the four vessel classes.
-Table 1: Class-wise configuration for the mixed statistical and structural textures dataset.
-Vessel Class Frequencies (Hz) Amplitude Distributions Modulation Rate
-[…]
-[…]
-[…]
-[…]
+where $w(t)=\exp\left(-\frac{(t-\mu)^{2}}{2\sigma^{2}}\right)$, with $\mu$ and $\sigma$ controlling the transition center and width.
+Each class employs distinct frequency bands, amplitude transition directions (e.g., K $\rightarrow$ Rayleigh), and modulation speeds.
+Lower-frequency classes (Vessel 1 and 2) are assigned slower modulation patterns to simulate large vessel behavior, while higher-frequency classes (Vessel 3 and 4) use faster modulations typical of smaller or more maneuverable ships [2].
 2.4 Statistical Texture Quantification via Bidirectional Temporal Entropy
-To quantify statistical texture in passive sonar signals, the bidirectional entropy-based method inspired by the temporal texturedness indicator proposed in [16] is adopted for this work.
-This approach evaluates how Shannon entropy accumulates over time in both forward and reverse directions of a signal.
-The underlying premise is that statistically textured signals exhibit consistent randomness over time, leading to minimal asymmetry between the two entropy trajectories.
-Let $x(t)$ be the input signal defined over time $t\in[0,T_{\text{obs}}]$, where $T_{\text{obs}}$ is the total observation duration.
-The signal is sampled at a frequency $f_{s}$ and divided into overlapping frames of length $L$ with hop size $H$.
-For each frame index $i$, the signal’s amplitude values are binned into a histogram with $M$ total bins.
-The resulting histogram is then normalized to obtain the probability distribution $\{p_{i,k}\}$, where $p_{i,k}$ denotes the probability corresponding to the $k$-th bin.
-The entropy at each step is computed using 6:
+Statistical texture is quantified using a bidirectional entropy-based approach inspired by [5], which evaluates how Shannon entropy accumulates over time in both forward and reverse directions.
+Let $x(t)$ denote the signal over $t\in[0,T_{\text{obs}}]$, sampled at rate $f_{s}$, and segmented into overlapping frames of length $L$ with hop size $H$.
+For each frame $i$, a normalized histogram with $M$ bins yields a probability distribution $\{p_{i,k}\}$, and entropy is computed as:
 $H_{i}=-\sum_{k=1}^{M}p_{i,k}\log_{2}p_{i,k}.$ (6)
-Entropy is accumulated in two directions: forward using the original signal $x(t)$, and reverse using the time-inverted signal $x_{r}(t)=x(T_{\text{obs}}-t)$.
-This results in two cumulative entropy curves for the forward and reverse directions, denoted $H_{d}(i)$ and $H_{r}(i)$, respectively.
-The surface area $S$ between these curves captures directional entropy asymmetry and is normalized by the theoretical maximum $S_{\text{max}}=T_{\text{obs}}\cdot\log_{2}M$, which corresponds to the entropy of a white noise signal.
-The relative entropy is defined as $P_{\text{rel}}=\frac{H_{\text{final}}}{\log_{2}M}$, where $H_{\text{final}}$ is the final cumulative entropy in the forward direction.
-The variable $P_{T}$ represents the proportion of the signal duration required for the forward and reverse entropy curves to converge.
-Using these components, the final Statistical Texture Score (StaTS) is defined in 7:
-$\text{StaTS}_{\text{stat}}=5\cdot P_{\text{rel}}\cdot\left(1-P_{T}\cdot\frac{S}{S_{\text{max}}}\right),$ (7)
-Higher scores reflect signals with strong statistical texture, characterized by high entropy, rapid forward–reverse convergence, and minimal directional asymmetry.
+Entropy is accumulated for the original signal $x(t)$ (forward) and its time-reversed form $x_{r}(t)=x(T_{\text{obs}}-t)$ (reverse), producing curves $H_{d}(i)$ and $H_{r}(i)$.
+The area $S$ between these captures directional asymmetry and is normalized by $S_{\text{max}}=T_{\text{obs}}\cdot\log_{2}M$.
+Relative entropy $P_{\text{rel}}=H_{\text{final}}/\log_{2}M$ and convergence proportion $P_{T}$ are also computed.
+The final statistical texture score is:
+$\text{StaTS}_{\text{stat}}=5\cdot P_{\text{rel}}\cdot\left(1-P_{T}\cdot\frac{S}{S_{\text{max}}}\right).$ (7)
+Higher scores indicate stronger statistical texture, marked by high entropy, fast convergence, and minimal asymmetry.
+This metric provides a compact and interpretable score that captures the degree of statistical texture by combining entropy magnitude, temporal symmetry, and convergence behavior.
 2.5 Structural Texture Quantification via Autocorrelation
-The Structural Texture Score (StrTS) is introduced to quantify periodicity and structural regularity in passive sonar signals using the normalized autocorrelation function.
-Autocorrelation measures the similarity between a signal and its time-shifted versions, making it effective for identifying repeating harmonic or modulated patterns [17].
-The signal is first transformed to be zero-mean before computing the autocorrelation.
-To ensure robustness across signals of different durations, the maximum lag is set adaptively as 25% of the total signal duration, with safeguards to prevent it from being too short or excessively long.
-This approach ensures that the method captures relevant structural patterns without being affected by noise or irrelevant long-term dependencies.
-Only positive lags beyond zero are considered, and the first significant positive peak is selected as the dominant periodic component.
-Negative peaks, indicating anti-phase alignment, are excluded.
-Let $R(\tau)$ denote the normalized autocorrelation value at lag $\tau$, where $\tau\in[1,\tau_{\text{max}}]$, and $\tau_{\text{max}}$ is the maximum lag considered.
-The lag corresponding to the first significant positive peak is identified as the dominant periodic component as shown in 8:
+Figure 2: Self-similarity matrices for three representative signals.
+(a) White noise (StrTS = 0.05): lacks clear structure.
+(b) Cargo ship signal (StrTS = 2.73): shows partial diagonal patterns.
+(c) Pure sine wave (StrTS = 5.00): exhibits strong and evenly spaced diagonals.
+Structural texture is measured using a normalized autocorrelation-based score, capturing periodicity and harmonic regularity [12].
+The signal is first zero-meaned, and the autocorrelation is computed up to a lag $\tau_{\text{max}}=0.25\cdot T_{\text{obs}}$.
+Only positive lags are considered, and the first significant peak is identified:
 $\tau^{*}=\arg\max_{\tau\in[1,\tau_{\text{max}}]}R(\tau),$ (8)
@@
 In addition to the scalar score, self-similarity matrices offer a complementary visual representation of structural repetition.
-These matrices capture pairwise similarity between overlapping frames of the signal, highlighting the repeated temporal pattern.
-Strong diagonal structures in the matrix correspond to high self-similarity, while noisy or unstructured matrices indicate a weak or lack of repetition.
-Figure 2: Self-similarity matrices for three representative signals.
-(a) White noise (StrTS = 0.05): lacks clear structure.
-(b) Cargo ship signal (StrTS = 2.73): shows partial diagonal patterns.
-(c) Pure sine wave (StrTS = 5.00): exhibits strong and evenly spaced diagonals.
-As shown in Fig.
-2, these visualizations qualitatively reflect periodic structure and align well with the StrTS scores.
-Higher StrTS scores correspond to clearer, more structured diagonal patterns, reinforcing the metric’s interpretability and relevance as a measure of structural texture.
+Clear, diagonal patterns correspond to higher StrTS scores.
+Figure 2 illustrates this with representative examples.
+The resulting score reflects structural periodicity in a normalized scale, making it useful for comparing a wide range of sonar signals and analyzing temporal regularity.
 3 Experimental Procedure
 3.1 Dataset Description
-Table 2: Dataset summary with number of classes and sample counts.
+Table 2: Summary of datasets with the number of classes and total sample counts.
 Dataset Type Classes Total Samples
@@
 The real-world datasets DeepShip [9] and Vessel Type Underwater Acoustic Data (VTUAD) [4] offer diverse passive sonar recordings.
-VTUAD is the largest with 175,965 samples across five classes.
-DeepShip provides 33,770 labeled samples across four classes.
-Table 2 summarizes the datasets by type, number of classes, and sample count.
+VTUAD is the largest with 175,965 samples across five classes, while DeepShip provides 33,770 labeled samples across four classes.
+Table 2 summarizes the datasets by type, number of classes, and total sample count.
 3.2 Experimental Design
-Table 3: Comparison of structural (StrTS) and statistical (StaTS) texture scores along with classification accuracies of TDNN[19] and HLTDNN [13] across five datasets.
-The last column shows the absolute accuracy improvement ($\Delta$ Acc) of HLTDNN over TDNN.
-The best average accuracy is bolded.
-Dataset StrTS StaTS TDNN Acc. (%) HLTDNN Acc. (%) $\Delta$ Acc. (%)
+[…]
+[…]
 Synthetic-Structural 4.96 $\pm$ 0.03 3.52 $\pm$ 0.09 72.50 $\pm$ 3.38 99.67 $\pm$ 0.15 +27.17
 Synthetic-Statistical 3.76 $\pm$ 0.35 4.93 $\pm$ 0.04 34.85 $\pm$ 1.04 55.05 $\pm$ 1.13 +20.20
-Synthetic-Mixed $4.39\pm 0.21$ $4.33\pm 0.56$ $72.18\pm 3.21$ $99.08\pm 0.88$ +26.90
+Synthetic-Balanced 4.39 $\pm$ 0.21 4.33 $\pm$ 0.56 72.18 $\pm$ 3.21 99.08 $\pm$ 0.88 +26.90
 DeepShip 3.16 $\pm$ 0.64 4.43 $\pm$ 0.55 50.49 $\pm$ 0.74 59.62 $\pm$ 1.69 +8.06
 VTUAD 4.45 $\pm$ 0.16 4.76 $\pm$ 0.32 71.92 $\pm$ 2.86 83.45 $\pm$ 0.57 +10.53
-All audio signals are resampled to 32 kHz for consistency.
-Each resampled signal is segmented into five-second intervals, except for the VTUAD dataset, which provides pre-segmented one-second clips with predefined dataset partitions.
-Segments are converted into logarithmic mel-frequency spectrograms using a Hanning window (window size 1024, hop 320) with 1024 mel-frequency bins.
-No data augmentation is applied to preserve the native texture characteristics.
-For all datasets (except VTUAD), samples are split into 70% training, 10% validation, and 20% test sets.
-Classification experiments use two models: TDNN as the baseline, and HLTDNN, which incorporates a histogram-based feature layer to enhance statisical texture representation.
+All signals are resampled to 32 kHz and segmented into five-second intervals, except VTUAD, which provides one-second clips with predefined train/validation/test splits.
+Log-mel spectrograms are computed using a Hanning window (size 1024, hop 320) with 1024 mel bins.
+No data augmentation is applied to preserve native texture characteristics.For all other datasets, a 70/10/20 train/validation/test split is used.
+TDNN serves as the baseline model, while HLTDNN incorporates a histogram-based feature layer to enhance statistical texture modeling.
 Both models are trained for 100 epochs using the Adagrad optimizer (learning rate $1\times 10^{-3}$, batch size 128).
-Test set accuracy is used for model evaluation.
-In addition to model performance, each dataset is analyzed using the two texture metrics (StrTS and StaTS) to provide insight into the texture properties of the signals.
+In addition to classification performance, each dataset is analyzed using the proposed texture quantification metrics: Structural Texture Score (StrTS) and Statistical Texture Score (StaTS).
+These scores provide insight into the texture proper- ties of each dataset and their influence on model performance.
 4 Results and Discussion
@@
 Table 3 summarizes the structural (StrTS) and statistical (StaTS) texture scores, along with classification accuracies for TDNN and HLTDNN across five datasets.
-The last column reports the accuracy improvement ($\Delta$ Accuracy) of HLTDNN over TDNN.
+The final column reports the absolute accuracy improvement ($\Delta$ Accuracy) of HLTDNN over TDNN.
 The results support the hypothesis that HLTDNN, with its histogram-based representation, is better suited for statistically rich signals.
-In the statistical texture dataset, where StaTS is highest ($4.93\pm 0.04$), HLTDNN achieves a substantial +20.20% improvement over TDNN.
-This validates HLTDNN’s capacity to capture statistical variations more effectively than TDNN.
-The structural texture dataset exhibits the highest StrTS ($4.96\pm 0.03$) and a lower StaTS.
-While TDNN performs reasonably well ($72.50\pm 3.38$), HLTDNN still outperforms it with a +27.17% gain, demonstrating that its statistical modeling does not hinder its ability to capture structural patterns.
-In the mixed statistical and structural textures dataset, where both texture types coexist, HLTDNN achieves near-perfect accuracy ($99.08\pm 0.88$) and shows a +26.90% gain over TDNN.
-This highlights HLTDNN’s adaptability in handling hybrid texture profiles.
-On the DeepShip dataset, which exhibits moderate statistical and lower structural texture complexity, HLTDNN outperforms TDNN by +8.06%.
-Despite being a real-world dataset with inherent variability, the gain shows HLTDNN’s effectiveness in generalizing to practical scenarios.
-VTUAD, the largest dataset, contains high values in both StrTS and StatS, and HLTDNN again shows a strong improvement of +10.53%.
-This result confirms the model’s scalability and ability to learn from complex, large-scale real-world data.
+In the statistical texture dataset, where StaTS is highest ($4.93\pm 0.04$), HLTDNN achieves a substantial +20.20% improvement over TDNN, validating its ability to capture statistical variations more effectively.
+The structural texture dataset exhibits the highest StrTS ($4.96\pm 0.03$) and lower StaTS.
+While TDNN performs reasonably well ($72.50\pm 3.38$), HLTDNN still surpasses it with a +27.17% gain, indicating that statistical modeling does not hinder its ability to capture structural patterns.
+In the mixed statistical and structural texture dataset, HLTDNN achieves near-perfect accuracy ($99.08\pm 0.88$) and shows a +26.90% gain over TDNN, demonstrating adaptability in handling hybrid texture profiles.
+On the DeepShip dataset, which presents moderate statistical and lower structural texture complexity, HLTDNN achieves an +8.06% improvement.
+Despite inherent variability in this real-world dataset, the performance gain reflects HLTDNN’s effectiveness in generalizing to practical scenarios.
+For the VTUAD dataset, which is the largest and exhibits high values in both StrTS and StaTS, HLTDNN delivers a strong +10.53% improvement.
+This confirms the model’s scalability and its ability to learn from complex, large-scale real-world data.
 4.2 Pretraining Models Using Synthetic Datasets
-Table 4: Model performance (%) on DeepShip with and without synthetic mixed texture pretraining. The best average accuracy is bolded.
-Model Without Pretraining With Pretraining
-TDNN $50.49\pm 0.74$ $52.5\pm 1.29$
-HLTDNN $59.62\pm 1.69$ $62.11\pm 1.36$
-To further investigate texture-aware learning, both models were pretrained on the mixed statistical and structural textures dataset and fine-tuned on DeepShip since the synthetic dataset was modeled after DeepShip’s characteristics.
-The performance comparison is shown in Table 4.
-Pretraining led to a consistent gain for both models, with HLTDNN improving by +2.49%.
-These results suggest that synthetic pretraining is particularly effective when the texture characteristics of the synthetic and real-world domains are well aligned.
-In summary, HLTDNN consistently outperforms TDNN across all datasets.
-The gains are especially notable in statistically textured conditions and further enhanced by texture-aligned pretraining.
-These findings highlight the benefit of explicitly modeling statistical texture and tailoring synthetic data to match real-world texture profiles.
+Table 4: Performance (%) on DeepShip with pretraining using different synthetic datasets. Gains are reported relative to the baseline: TDNN = 50.49 ± 0.74, HLTDNN = 59.62 ± 1.69.
+Pretraining Dataset TDNN HLTDNN
+Accuracy Gain Accuracy Gain
+Synthetic-Statistical 47.21 ± 1.89 –3.28 53.21 ± 1.89 –6.41
+Synthetic-Structural 50.63 ± 1.95 +0.14 56.88 ± 1.22 –2.74
+Synthetic Statistical and Structural 52.50 ± 1.29 +2.01 62.11 ± 1.36 +2.49
+To evaluate the impact of synthetic pretraining, both models were pretrained on synthetic datasets and then fine-tuned on DeepShip.
+As shown in Table 4, pretraining with the mixed texture dataset improved performance for both TDNN and HLTDNN, with a gain of +2.01% and +2.49%, respectively.
+In contrast, statistical-only or structural-only pretraining resulted in minimal or negative gains.
+This indicates that synthetic pretraining is particularly effective when the texture characteristics of the synthetic and real-world domains are well aligned.
+The mixed synthetic statistical & structural dataset, which captures both statistical and structural properties, more closely reflects the texture complexity present in DeepShip, leading to better generalization after pretraining.
 5 Conclusion
-This work investigated the role of structural and statistical texture in CNN-based passive sonar classification.
-Statistically rich signals frequently occur in passive sonar environments due to stochastic propagation, scattering, and background interference — phenomena that standard CNNs may fail to capture effectively.
-By constructing synthetic underwater datasets with controlled texture attributes, it was possible to isolate and analyze model behavior across varying degrees of signal complexity.
-Experimental findings show that traditional CNN architectures such as TDNN are more sensitive to structural regularities, whereas the texture-aware HLTDNN demonstrates improved performance on signals characterized by strong statistical texture.
-These results support the importance of explicitly modeling statistical texture in passive acoustic target classification.
-The proposed statistical and structural texture quantification metrics further contribute to model interpretability by linking signal properties to performance characteristics.
-Future directions include generating synthetic data samples to adapt audio foundation models such as AudioPaLM [20] and AudioGPT [21] for passive sonar, analyzing texture sensitivity in transformer-based architectures such as the Audio Spectrogram Transformer (AST) [22], embedding texture-aware regularization into training objectives, leveraging more advanced approaches for data generation such as diffusion models [23] and applying the texture metrics to other audio domains.
-6 Acknowledgment
-Portions of this research were conducted with the advanced computing resources provided by Texas A&M High Performance Research Computing.
+This work explored the role of structural and statistical texture in CNN-based passive sonar classification.
+Statistically rich signals, common in passive sonar due to stochastic propagation and interference, are often underrepresented in standard CNN processing.
+Synthetic datasets with controlled texture characteristics enabled systematic evaluation of model performance under varying texture complexities.
+Results show that while TDNN favors structural regularities, the texture-aware HLTDNN performs better on statistically textured signals, highlighting the value of modeling statistical texture.
+The proposed texture quantification metrics offer interpretability by linking signal properties to model behavior.
+Future work includes adapting foundation models like AudioPaLM [20] and AudioGPT [8], evaluating transformer-based models such as AST [6], integrating texture-aware regularization, using diffusion-based generation [11], and extending this framework to other audio domains.